A fuel tank holds 22.3 gallons of gasoline. If the density of gasoline is 0.8206 g/mL, what is the mass in kilograms of gasoline in a full tank?
1
Expert's answer
2019-09-05T03:33:12-0400

Since

"Density = Mass\/Volume"


1 Gallon = 3.785 L = 3785 ml

22 Gallons = 3.785 * 22 = 83.27 L = 83270 ml

Density = 0.8206 g/ml


0.8206 = m/83270

m = 68331.36 g = 68.33 Kg